# Accademia dei Lincei, Roma

### v1

@@ 3 Weak-discontinuity waves
For system (1), the theory of weak discontinuity (acceleration) waves is classical and will not be recalled here; we refer to Boillat for the three-dimensional case Boillat and to Ruggeri ; Newbook for full details in one space dimension.
@@
$a=(\nabla\lambda\cdot\mathbf{d})_{E},\qquad b=-(\mathbf{l}\cdot\nabla\mathbf{f}\cdot\mathbf{d})_{E},$
@@ 4 Nonlinear model for uniaxial isothermal viscoelasticity and non-Newtonian fluids
$P(F,0)=0,\qquad P_{F}(F,0)=0,\qquad P_{\sigma}(F,0)\leq 0.$
@@ 5 Acceleration waves for the present model
$\mathbf{u}\equiv(v,F,\sigma)^{\top},\quad\mathbf{G}\equiv\left(-\frac{T(F)+\sigma}{\rho^{*}},-v,-\frac{v}{\omega(\sigma)}\right)^{\top},\quad\mathbf{f}\equiv\left(0,0,\frac{P(F,\sigma)}{\omega(\sigma)}\right)^{\top}.$
Accordingly,
$\mathbf{A}=\nabla\mathbf{G}\equiv\left(\begin{array}[]{ccc}0&-\dfrac{{\cal{W}}^{{}^{\prime\prime}}(F)}{\rho^{*}}&-\dfrac{1}{\rho^{*}}\\ -1&0&0\\ -\dfrac{1}{\omega(\sigma)}&0&0\end{array}\right).$
@@
$\bm{\nabla}\mathbf{f}=\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&\dfrac{P_{F}(F,\sigma)}{\omega(\sigma)}&\dfrac{P_{\sigma}(F,\sigma)\omega(\sigma)-P(F,\sigma)\omega^{\prime}(\sigma)}{\omega^{2}(\sigma)}\end{array}\right).$
Assuming that the wave propagates in an undeformed equilibrium state, $F=1$ and $\sigma=0$, and using (11), we obtain
$a=\frac{-\omega^{\prime}(0)+\omega(0)^{3}{\cal{W}}^{{}^{\prime\prime\prime}}(1)}{2\lambda_{0}^{3}\rho^{*}\omega(0)^{3}},\qquad b=-\frac{P_{\sigma}(1,0)}{2\rho^{*}\lambda_{0}^{2}\omega^{2}(0)},$ (13)
@@
$t_{c}=-\frac{1}{b}\ln\!\left(1-\frac{{\cal{G}}_{\rm cr}}{{\cal{G}}_{0}}\right),\qquad{\cal{G}}_{0}>{\cal{G}}_{\rm cr}.$ (15)
@@
Case 3: $P_{\sigma}(1,0)\sim\varepsilon^{-n}$, $n>1$, where $\varepsilon$ is a small parameter (thus $b(\varepsilon)\to+\infty$ as $\varepsilon\to 0$).
@@ 6 Acceleration waves in viscoelastic solids under uniaxial deformation
However, in order to compute explicitly $\lambda_{0}$ and hence ${\cal{G}}_{\rm cr}$, we also need $\omega^{\prime}(0)$ and ${\cal{W}}^{{}^{\prime\prime\prime}}(1)$, which requires a third-order approximation of the potentials.
@@
For the elastic potential, the precise expression depends on the chosen constitutive model. Nevertheless, since $W$ and all its derivatives have the same physical dimension as $E_{1}$, we may write, near equilibrium,
@@
$\omega^{\prime}(0)=0,\qquad{\cal{W}}^{{}^{\prime\prime\prime}}(1)=-E_{1}R<0.$
@@
$\lambda_{0}=\sqrt{\frac{E_{1}+E_{2}}{\rho^{*}}},\quad a=-\frac{\sqrt{\rho^{*}}\,E_{1}R}{2\,(E_{1}+E_{2})^{3/2}}<0,\quad b=\frac{E_{2}}{2\tau(E_{1}+E_{2})}>0,$ (21)
where the relaxation time $\tau$ is related to the viscosity coefficient $\mu_{0}$ by (see (19))
$\tau=\frac{\mu_{0}}{E_{2}}.$
@@
${\cal{G}}_{\rm cr}=\frac{b}{|a|}=\frac{1}{R\tau}\frac{E_{2}}{E_{1}}\sqrt{\frac{E_{1}+E_{2}}{\rho^{*}}}.$ (22)
@@ 6.1 Evaluation of $R$ with the Mooney–Rivlin potential
${\cal W}^{\prime\prime}(1)=E_{1}=2.12~\mathrm{MPa},\qquad{\cal W}^{\prime\prime\prime}(1)=-E_{1}R=-6.93~\mathrm{MPa},$
@@
$E_{2}=3~\mathrm{MPa},\qquad\tau=0.1~\mathrm{s},\qquad\rho^{*}=929~\mathrm{kg/m^{3}}.$
@@
$\lambda_{0}\approx 74.24~\mathrm{m/s},\quad a\approx-0.009~\mathrm{s/m},\quad b\approx 2.93~\mathrm{s^{-1}},\quad{\cal{G}}_{\rm cr}\approx 321.27~\mathrm{m/s^{2}}\approx 32.79\,g.$
@@ 7 Acceleration waves in a non-Newtonian fluid
$\bm{\sigma}=k\,\dot{\gamma}^{\,m-1}\mathbf{D},\qquad\dot{\gamma}=2|\mathbf{D}|,\qquad|\mathbf{D}|=\sqrt{\mathbf{D}:\mathbf{D}},$ (24)
@@
We now consider, also in this case, a quadratic viscous energy (18)2, for which $\omega$ is constant and given by
$\omega=\frac{\mu_{0}}{\tau_{0}}.$ (28)

### v2

@@ Abstract
dedicatory: With great esteem and fraternal affection, this paper is dedicated to Marco Sammartino on the occasion of his 60th birthday.
@@ 3 Weak-discontinuity waves
The theory of weak discontinuity (acceleration) waves is classical and will not be recalled here; we refer to Boillat for the three-dimensional case Boillat and to Ruggeri ; Newbook for full details in one space dimension.
@@
$a=(\nabla\lambda\cdot\mathbf{d})_{E},\qquad b=-(\mathbf{l}\cdot\nabla\mathbf{f}\cdot\mathbf{d})_{E},$ (6)
@@ 4 Nonlinear model for uniaxial isothermal viscoelasticity and non-Newtonian fluids
$P(F,0)=0,\qquad P_{F}(F,0)=0,\qquad P_{\sigma}(F,0)\leq 0.$ (13)
@@ 5 Acceleration waves for the present model
$\mathbf{u}\equiv(\rho^{*}v,F,Z(\sigma))^{\top},\quad\mathbf{G}\equiv\left(-\bigl(T(F)+\sigma\bigr),-v,-v\right)^{\top},\quad\mathbf{f}\equiv\left(0,0,P(F,\sigma)\right)^{\top}.$
As for acceleration waves, the balance form is not needed. It is therefore convenient to rewrite system (1) in the equivalent quasi-linear form
$\partial_{t}\mathbf{u}+\mathbf{A}(\mathbf{u})\,\partial_{X}\mathbf{u}=\mathbf{f}(\mathbf{u}),$ (14)
by substituting the field and production terms as
$\mathbf{u}\equiv(v,F,\sigma)^{\top},\qquad\mathbf{f}\equiv\left(0,0,\frac{P(F,\sigma)}{\omega(\sigma)}\right)^{\top}.$ (15)
With this choice of the field (15)1, the characteristic matrix (14) associated with the system (9) becomes
$\mathbf{A}\equiv\left(\begin{array}[]{ccc}0&-\dfrac{{\cal W}^{\prime\prime}(F)}{\rho^{*}}&-\dfrac{1}{\rho^{*}}\\ -1&0&0\\ -\dfrac{1}{\omega(\sigma)}&0&0\end{array}\right).$
@@
$\bm{\nabla}\mathbf{f}=\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&\dfrac{P_{F}(F,\sigma)}{\omega(\sigma)}&\dfrac{P_{\sigma}(F,\sigma)\omega(\sigma)-P(F,\sigma)\omega^{\prime}(\sigma)}{\omega(\sigma)^{2}}\end{array}\right).$
Assuming that the wave propagates in an undeformed equilibrium state, $F=1$ and $\sigma=0$, and using (13) and (6) we obtain
$a=\frac{-\omega^{\prime}(0)+\omega(0)^{3}{\cal{W}}^{{}^{\prime\prime\prime}}(1)}{2\lambda_{0}^{3}\rho^{*}\omega(0)^{3}},\qquad b=-\frac{P_{\sigma}(1,0)}{2\rho^{*}\lambda_{0}^{2}\omega(0)^{2}},$ (17)
@@
$t_{c}=-\frac{1}{b}\ln\!\left(1-\frac{{\cal{G}}_{\rm cr}}{{\cal{G}}_{0}}\right).$ (19)
@@
Case 3: $P_{\sigma}(1,0)\sim-\varepsilon^{-n}$, $n>1$, where $\varepsilon$ is a small parameter (thus $b(\varepsilon)\to+\infty$ as $\varepsilon\to 0$).
@@ 6 Acceleration waves in viscoelastic solids under uniaxial deformation
However, in order to compute explicitly $a$ and hence ${\cal{G}}_{\rm cr}$, we also need $\omega^{\prime}(0)$ and ${\cal{W}}^{{}^{\prime\prime\prime}}(1)$, which requires a third-order approximation of the potentials.
@@
For the elastic potential, the precise expression depends on the chosen constitutive model. Nevertheless, since $\cal{W}$} and all its derivatives have the same physical dimension as $E_{1}$, we may write, near equilibrium,
@@
$\omega^{\prime}(0)=0,\qquad{\cal{W}}^{{}^{\prime\prime\prime}}(1)=-2E_{1}R<0.$
@@
$\lambda_{0}=\sqrt{\frac{E_{1}+E_{2}}{\rho^{*}}},\quad a=-\frac{\sqrt{\rho^{*}}\,E_{1}R}{(E_{1}+E_{2})^{3/2}}<0,\quad b=\frac{E_{2}}{2\tau_{0}(E_{1}+E_{2})}>0,$ (25)
where the relaxation time $\tau_{0}$ is related to the viscosity coefficient $\mu_{0}$ by (see (23))
$\tau_{0}=\frac{\mu_{0}}{E_{2}}.$ (26)
@@
${\cal{G}}_{\rm cr}=\frac{b}{|a|}=\frac{1}{2R\tau_{0}}\frac{E_{2}}{E_{1}}\sqrt{\frac{E_{1}+E_{2}}{\rho^{*}}}.$ (27)
@@ 6.1 Evaluation of $R$ with the Mooney–Rivlin potential
${\cal W}^{\prime\prime}(1)=E_{1}=2.12~\mathrm{MPa},\qquad{\cal W}^{\prime\prime\prime}(1)=-2E_{1}R=-6.93~\mathrm{MPa},$
@@
$E_{2}=3~\mathrm{MPa},\qquad\tau_{0}=0.1~\mathrm{s},\qquad\rho^{*}=929~\mathrm{kg/m^{3}}.$
@@
$\lambda_{0}\approx 74.21~\mathrm{m/s},\quad a\approx-0.009~\mathrm{s/m},\quad b\approx 2.93~\mathrm{s^{-1}},\quad{\cal{G}}_{\rm cr}\approx 321.41~\mathrm{m/s^{2}}\approx 32.80\,g.$
@@ 7 Acceleration waves in a non-Newtonian fluid
$\bm{\sigma}=k\,\dot{\gamma}^{\,m-1}\mathbf{D},\qquad\dot{\gamma}=2\lVert\mathbf{D}\rVert,\qquad\lVert\mathbf{D}\rVert=\sqrt{\mathbf{D}:\mathbf{D}}$ (29)
@@
We now consider, also in this case, a quadratic viscous energy (22)2 with (26), for which $\omega$ is constant and given by
$\omega=\frac{\tau_{0}}{\mu_{0}}.$ (33)